OpenStudy (mathhelp346):

Write an equation and solve: Four less than three times a number is three more than two time the number. What is the number?

OpenStudy (mathhelp346):

I tried 3t-4=2t-3 but it didnt work

OpenStudy (tyteen4a03):

Let the number be x,the equation you're looking for is 4 - 3x = 3 + 2x.